please help solve the equation and graph on a number line. equation 3x+2=7x-6. this is how i did it. 3x=2-3x=7x-6-3x 2=4x-6 then i added 6 to 2 making 8 on the other side the 6 cancles out leaving 4x 4x divided by 4 leaves x on the other side 8 divided by 4 leaves 4 solution x=4 but when i check it the sides do not equal. thanks Sandra D.
Answer by wuwei96815(245)

3x+2=7x-6
3x-7x=-6-2
-4x=-8
x=2